# Underwater Image Enhancement by Diffusion Model with Customized CLIP-Classifier

### v1

@@ III-A Preliminary of Diffusion Models
where $\mu_{\theta}(x_{t},t)$ is a function of the mean, and $\sigma_{\theta}(x_{t},t)$ is a fuction of the variance. Some works [22, 45] set the variance of the two Gaussians of the forward and reverse processes to match exactly, and the diffusion model just need to optimize $\mu_{\theta}(x_{t},t)$. Then, when $x_{t}$ is available, $\mu_{\theta}(x_{t},t)$ can be reparameterized as:
@@
where $\epsilon_{\theta}(x_{t},t)$ is a neural network resembles a function approximator trained to predict $\epsilon$ from $x_{t}$, and and the optimizing objective of Evidence Lower Bound (ELBO) is converted from $\mu_{\theta}(x_{t},t)$ to $\epsilon_{\theta}(x_{t},t)$, as follows:
@@
For another understanding for the diffusion model, Song et al. [31, 46] unifies the diffusion model, a discrete multi-step denoising process, into a special form of a continuous stochastic differential equation (SDE), and the reverse process to simulate the score of each marginal distribution $\bigtriangledown\log p(x_{t})$ from $T$ to 00. Recall that $\bigtriangledown$ is shorthand for $\bigtriangledown_{x_{t}}$ in the interest of brevity. Meanwhile, the score matching of the score-based generative model is equivalent to ELBO optimization, and the score $\bigtriangledown\log p(x_{t})$ can be represented as the linear mapping of $\epsilon_{\theta}(x_{t},t)$ with the following relationship:
@@
For conditional generation models, Dhariwal et al. [24] uses gradients from a classifier $p(y|x_{t})$ to guide the reverse process, and by Bayes rule, the score function of this is formed as
@@
Notably, the gradient of $\bigtriangledown\log p(y)$ with respect to $x_{t}$ is zero, $\bigtriangledown\log p(x_{t})$ is defined as unconditional score, and $\bigtriangledown\log p(y|x_{t})$ is defined as adversarial gradient [45]. Then, $\epsilon_{\theta}(x_{t},t)$ is redefined as $\epsilon_{\theta}(x_{t},y,t)$ which corresponds to the score of the joint distribution [24]:
@@
Simultaneously, the optimizing objective is also rewritten as:
@@
III-B Pipeline of Underwater Enhancement
III-B1 Multi-Classifier for Diffusion Models
Our method leverages a diffusion model for underwater image enhancement, and we use the image-to-image model SR3 [25] as the base diffusion framework. To condition the SR3 on the input $y$, the input $y$ is concatenated with $x_{t}$ along the channel dimension (see in Fig. 2 (b)), and experiments have shown that simple concatenation yields similar generation quality and makes the model more robust [25].
For conditional diffusion models, the most popular form of guidance is known as Classifier Guidance [32, 31], which is derived from score-based generation models. Our goal is to train a neural network to predict $\bigtriangledown\log p(x_{t}|y)$ the score of the conditional model, at arbitrary transition steps. Recall the score function from Equation (7):
@@ III-B1 Multi-Classifier for Diffusion Models
where $y$ is an underwater image in this work, and of course, it can be other conditions like prompts, Canny edges, etc [21]. Following Song et al. [31] and Luo et al. [45], the Classifier Guidance can be interpreted as learning an unconditional score function combined with the adversarial gradient of a classifier $p(y|x_{t})$, and building on this observation, we will later expand the single classifier into multi-classifiers. In the following, we divide our approach into two stages: the stage-one is based on the single classifier to learn prior knowledge of the synthetic paired dataset, and the stage-two is based on the multi-classifier to fine-tune the model with additional dataset.
Stage-one: Single Classifier Guidance. Given the paired dataset synthesized by color transfer, denoted $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{N}$, and we are interesting in training the conditional diffusion model to learn the prior knowledge mapping from the real underwater degradation domain to the real in-air natural domain with the learning objective $\mathcal{L}_{2}$ (refer to Eq. (9)), over $x$ is an in-air natural image, where the condition $y$ is a synthesize degraded image (see in Fig. 2 (b)). Utilizing this prior knowledge can counteract the adverse impact of unreal synthetic. However, this model will be guided by the latter condition without real-underwater reference, and the fine-tuning, to directly continue training with the additional paired data, may destroy the prior knowledge.
To avoid this issue, thanks to the Classifier Guidance strategy, we first train a classifier with prior knowledge, distinguishing between in-air natural domain images and underwater domain images. Then, we exploit this classifier and the additional paired datasets to jointly fine-tune the diffusion model. As shown in Fig. 3, the enhancement results are guided by the combination of the reference domain and the in-air natural domain.
Stage-two: Multiple Classifiers Guidance. Using Bayes rule, the score function of the multi-condition model under Classifier Guidance can be derived as:
$\bigtriangledown\log p(x_{t}|y_{1},y_{2})$ $=\bigtriangledown\log(\frac{p(x_{t})p(y_{1},y_{2}|x_{t})}{p(y_{1},y_{2})})$ (10)
=$\bigtriangledown$= ▽ $\log p(x_{t})+\bigtriangledown\log p(y_{1},y_{2}|x_{t}).$
$\bigtriangledown\log p(y_{1},y_{2}|x_{t})=\bigtriangledown\log p(y_{1}|x_{t})+\bigtriangledown\log p(y_{2}|x_{t}).$ (11)
Then, substituting this into Equation (10), we have:
$+\bigtriangledown\log p(y_{2}|x_{t}),$
where $p(y_{1}|x_{t})$ is the same as $p(y|x_{t})$ in Eq. (7), and $p(y_{2}|x_{t})$ is the trained classifier. By simply recursion, it is possible to achieve multi-conditions guidance by simple adding conditional independent classifiers.
$+(1-\lambda)\bigtriangledown\log p(y_{2}|x_{t}),$
when $\lambda<0.5$, the conditional diffusion model prioritizes the conditioning information $y_{2}$ and moves more towards the in-air natural domain. The value of $\lambda$ is empirically set to 0.4 in our experiments. Meanwhile, according to Eq. (6) and Eq. (9), we rewrite $\epsilon_{\theta}(x_{t},t)$ as $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$, combining the adversarial gradient of the two classifiers:
and the overall learning objective $\mathcal{L}_{3}$ of the multi-classifiers diffusion model is depicted in Equation (15):
$\mathcal{L}_{3}=\min_{\theta}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),y_{1},y_{2},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},y_{1},y_{2},t)\right\|^{2},$ (15)
where this learning objective $\mathcal{L}_{3}$ is directly used in fine-tuning the pre-trained diffusion model with additional paired datasets.
Figure 3: New fine-tuning strategy. From state $x_{T}$, according to Eq. 9, we first adopt a single classifier guidance strategy, using condition $y_{1}$—the input source image—to guide the reverse diffusion process until state $x_{t}$. Then, we switch to the multi-classifier guidance strategy according to Eq. 15. With multi-condition guidance, the intermediate results from $x_{t}$ to $x_{0}$ are constrained by both the natural and reference domains, mitigating the damage of fine-tuning to the prior knowledge of the pre-trained model.
III-B2 Color Transfer for Degraded Image Synthesis
The color transfer can realize the translation between image domains, by transfering one image’s color characteristics to another [29]. Naturally, we can use this method to translate the image from the in-air natural domain into the underwater scene domain.
@@ III-B2 Color Transfer for Degraded Image Synthesis
In our case, for each in-air natural image $I_{s}$, we randomly select one image as Template A from Template Pool (the underwater scene domain) as the target image $I_{t}$, and then use $I_{t}$ to guide $I_{s}$ to transfer to the underwater scene domain via Equation (16). The detail process is depicted in Fig. 2 (a). Finally, we convert the degraded synthetic image back to RGB color space.
@@
Figure 4: Illustration of the prompt learning for CLIP-Classifier. (A) Prompt Initialization. Given two text prompts describing the in-air natural image $I_{n}$ and underwater image $I_{u}$. We encode each text and get the initial in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$ and the initial underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$. (B) Prompt Training. We use the cross-entropy loss to constrain the learnable prompts, aligning learnable prompts and images in the CLIP latent space, by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$. The base model of CLIP is frozen throughout. The learned prompt for CLIP-Classifier will be used in Fig. 3
III-B3 Prompt Learning for CLIP-Classifier
@@
In Section. III-B1, what we want is a classifier to control the diffusion model generation process, making the result move toward the in-air natural domain. Thanks to the CLIP’s near-perfect text-driven property [43], we design a CLIP-Classifier to implement this vision, and all we need to do is to find prompts that can distinguish between in-air natural images and underwater images. However, prompts for the quality and abstract perception of natural and underwater images often require a lot of manual prompt tuning, and the similar prompts easily lead to huge differences in CLIP score. To overcome this issue, CoOp [28] models a prompt’s context words with learnable vectors to broader CLIP-like vision-language models to downstream applications. For our task, we only need two prompts to characterize in-air natural and underwater images. Therefore, similar to CLIP-LIT [27], we utilize the prompt learning to model a prompt’s the longest text with learnable tensors, and maximize the ability to characterize images. Note that the base model of CLIP is frozen in the entire process. The process of prompt learning involves two key components: prompt initialization and prompt training, as detailed below.
@@
Prompt Initialization. Randomly given two texts describing in-air natural and underwater images respectively, CLIP converts each text into the multi-model embedding space, obtaining the in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$, and the underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$ (refer to Fig. 4 (a)). $N$ represents the length of prompt tokens, with an upper limit of 77. CoOp [28] has experimentally shown that more prompt tokens lead to better performance. Thus, we model $T_{n}$ and $T_{u}$ as learnable tensors with $N=77$, maximizing the length of prompt tokens.
@@
Prompt Training. Given an in-air natural image $I_{n}$ and an underwater image $I_{u}$, the image encoder $\Phi(\cdot)$ of the frozen CLIP takes them as input and extracts image features. On the other hand, by forwarding the prompt $T_{n}$ and the underwater prompt $T_{u}$ to the text encoder $\Psi(\cdot)$, we can obtain the text features. Given a batch of image-text pairs, the prediction probability is calculated as follows:
@@
where $I$ is an image, $I\in\left\{I_{n},I_{u}\right\}$ and $cos(\cdot,\cdot)$ denotes cosine similarity. Specially, the $T_{n}/T_{u}$ is shared among all in-air natural/underwater images. Similar to CLIP [43], the prompt learning is trained to align learnable prompts and images in embedding spaces, by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$, while minimizing the cosine similarity for incorrect pairs. The cross-entropy loss can be used as a learning objective, and its gradients can be propagated all the way back to update the prompt tensors (see in Fig. 4 (b)). Since there are only two prompts in our prompt engineering, we can simplify the process by directly using the binary cross-entropy to classify in-air natural and underwater images to learn the prompt tensors:
@@
Substituting CLIP-Classifier $\mathcal{L}_{clip}$ into Equation (14), the new $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$ is:
@@
Where due to the properties of the cosine function, $y_{2}=T_{u}$. Next, we explore the details of the CLIP-Classifier $\mathcal{L}_{clip}$ guides generation process.
@@
III-C Fine-Tuning Strategy
The forward process of the diffusion model is actually a process in which image semantic information is continuously obscured by noise, always starting from high-frequency information to low-frequency information. In contrast, the reverse generation process smoothly converts an initial Gaussian noise to a realistic image sample through iterative denoising, working from low-frequency information to high-frequency information. For special image-to-image tasks that require consistent content, such as super-resolution [25], image restoration [47], repaint [48], etc., the input image and the corresponding image share low-frequency content in each transition from $x_{t}$ to $x_{t-1}$ during the generation process [49]. Therefore, image synthesis and editing on diffusion model can start from the high-frequency region. For instance, SDEdit [50] is to “hijack” the generation process, adding a suitable amount of noise to smooth out high-frequency semantic details, while still preserving the overall structure of the input image, and then initialize the SDE with this noisy input, obtaining a denoised result that is faithful to the user guidance input.
Obviously, in our case, the underwater degradation images and the reference images share low-frequency content, so we can introduce new conditions at intermediate moments $x_{t}$ to guide the diffusion process (see in Fig. 3). In addition, given a batch of data pairs, we first use Eq. (2) to produce a sequence of intermediate variables $(x_{0},x_{1},\dots,x_{T})$ for each image, by perturbing images with Gaussian noise at time $t\in(0,1)$, and then employ CLIP-Classifier to calculate the classification probability (CLIP score), as illustrated in Fig. 5 (a). Notably, the randomness of the perturbation noise directly causes the CLIP score curves to vibrate violently in small intervals, but the overall trend of the curves is still pronounced. From Fig. 5 (b), we observe that the CLIP-Classifier $\mathcal{L}_{clip}$ mainly works in the discontinuous high-frequency region where $t\in(0,0.10)\cup(0.25,0.50)$, and other range can not be distinguished by the CLIP-Classifier. Combined with the diffusion model generative prior, sharing the low-frequency content, we only need to guide the pre-trained diffusion model in the high-frequency region during the fine-tuning process. The new strategy improves the fine-tuning speed compared to the traditional strategy, and can even speed up to 10 times, as will be discussed in detail in Section IV-F.

### v2

@@ III-A Preliminary Knowledge of Diffusion Models
where $\mu_{\theta}(x_{t},t)$ and $\sigma_{\theta}(x_{t},t)$ are the functions of mean and variance, respectively. Some works [22, 45] set the two Gaussian variances of the forward and reverse processes to match exactly, and the diffusion model just need to optimize $\mu_{\theta}(x_{t},t)$. Then, when $x_{t}$ is available, $\mu_{\theta}(x_{t},t)$ can be reparameterized as:
@@
where $\epsilon_{\theta}(x_{t},t)$ is a neural network similar to a function approximator trained to predict $\epsilon$ from $x_{t}$, and the optimizing objective of Evidence Lower Bound (ELBO) is converted from $\mu_{\theta}(x_{t},t)$ to $\epsilon_{\theta}(x_{t},t)$, as follows:
@@
As another understanding of the diffusion model, Song et al. [31, 46] unify the diffusion model, a discrete multi-step denoising process, into a special form of a continuous stochastic differential equation (SDE), and the reverse process is to simulate the score of each marginal distribution $\bigtriangledown\log p(x_{t})$ from $T$ to 00. Recall that $\bigtriangledown$ is shorthand for $\bigtriangledown_{x_{t}}$ in the interest of brevity. Meanwhile, the score matching of the score-based generative model is equivalent to ELBO optimization, and the score $\bigtriangledown\log p(x_{t})$ can be represented as the linear mapping of $\epsilon_{\theta}(x_{t},t)$ with the following relationship:
@@
Figure 2: The preparation for the pre-trained model. (a) Randomly select a Template from the Template Pool (underwater domain). Then, the Color Transfer module, guided by the template, degrades an in-air natural image from INaturalist [34] into underwater domain, constructing paired datasets for training image-to-image diffusion model. (b) The image-to-image diffusion model SR3 [25] is trained to learn the prior knowledge of mapping transitions from the real underwater degradation domain to the real in-air natural domain and to generate the corresponding enhancement results based on input synthetic underwater images produced by Color Transfer.
For conditional generation models, Dhariwal et al. [24] uses gradients from a classifier $p(y|x_{t})$ to guide the reverse process, where $y$ is the condition information. By Bayes formula, the score function of this is formed as
@@
Notably, the gradient of $\bigtriangledown\log p(y)$ with respect to $x_{t}$ is zero, $\bigtriangledown\log p(x_{t})$ is defined as an unconditional score, and $\bigtriangledown\log p(y|x_{t})$ is defined as the adversarial gradient [45]. Then, $\epsilon_{\theta}(x_{t},t)$ is redefined as $\epsilon_{\theta}(x_{t},y,t)$ which corresponds to the score of the joint distribution [24]:
@@
Simultaneously, the optimizing objective is rewritten as:
@@
III-B Pipeline of Underwater Image Enhancement
III-B1 Multi-Guidance for Diffusion Models
Our method leverages a diffusion model for underwater image enhancement, and we use the image-to-image model SR3 [25] as the base diffusion model framework. To condition the SR3 on the input underwater image $y$, the input $y$ is concatenated with the intermediate variable $x_{t}$ along the channel dimension (see in Fig. 2 (b)), and experiments have shown that simple concatenation produces similar generation quality and enhances the model’s robustness [25].
The most prevalent form of guidance for conditional diffusion models is Classifier Guidance [32, 31], which is derived from score-based generation models. Our goal is to train a neural network to predict $\bigtriangledown\log p(x_{t}|y)$, the score of the conditional model, at arbitrary transition steps. Recall the score function from Eq. (7):
@@ III-B1 Multi-Guidance for Diffusion Models
where $y$ represents the conditional information, which is an underwater image in this work, and of course, it can be other conditions like prompts, Canny edges, human poses, etc [21]. Additionally, a classifier $p(y|x_{t})$ is used to take in arbitrary noisy $x_{t}$ and attempt to predict conditional information $y$. Directly using the gradients $\bigtriangledown\log p(x_{t}|y)$ constructed from the underwater image $y$ and the manually selected reference image $x_{0}$ to control SR3 may produce competitive enhancement results, but its upper performance bound is limited by the reference domain. The common approach to address this problem is to leverage the prior knowledge from pre-trained models to counteract the limitation of the reference domain. However, directly fine-tuning diffusion models with UIE benchmark datasets tends to compromise this prior knowledge and makes it difficult to fully exploit the diffusion models’ potential.
In Classifier Guidance [31, 45], the score $\bigtriangledown\log p(x_{t}|y)$ can be interpreted as learning an unconditional score function $\bigtriangledown\log p(x_{t})$ combined with the adversarial gradients of a classifier $p(y|x_{t})$. Building on this observation, we can easily combine Bayes formula to expand a single classifier into multiple classifiers and achieve multi-guidance of diffusion models. This discover offers a fresh perspective, enabling us to introduce more prior knowledge of different domains in the form of classifiers to jointly fine tune the diffusion model, thereby breaking the limitation of the reference domain. Specially, we divide our approach, abbreviated as CLIP-UIE, into two stages. In the first stage, we use the synthetic dataset produced by color transfer as the single conditional information to guide the image-to-image diffusion model SR3. This approach aims to enable SR3 to capture the prior knowledge of mapping transitions from the underwater degradation domain to in-air natural domain. In the second stage, we exploit a trained classifier that possesses the prior knowledge of the in-air natural domain. Then we combine it with the UIE benchmark dataset to jointly fine-tune the pre-trained baseline SR3, guiding the enhancement results towards the in-air natural domain and bridging the domain gap between synthetic and real images.
Stage-one: Obtain the Pre-trained Model with Prior Knowledge of Domain Adaption.
(a) Data Preparation: Natural images can be transformed into underwater images through a light scattering model [20]. However, the degradation type of this kind of synthetic images is relatively simple and limited in number, which cannot simulate diverse underwater scenes. Considering that color transfer can efficiently and intuitively translate between image domains by transferring the color characteristics of one image to another [29], it enlightens and enables us to transform natural images captured in air into corresponding degraded underwater images. The more images of different underwater degradation types we provide, the closer the synthetic dataset gets to the real underwater domain.
@@
In our case, for each in-air natural image $I_{s}$, we randomly select one image from Template Pool (the underwater scene domain) to serve as the target image $I_{t}$, which also functions as Template. We then use $I_{t}$ to guide $I_{s}$ in transitioning to the underwater scene domain via Eq. (10). The detailed process is depicted in Fig. 2 (a). Finally, we convert the degraded synthetic image back to RGB color space. To facilitate the distinction, our synthetic dataset created by color transfer is referred to as the UIE-air dataset, while the synthetic dataset obtained by manual selection is referred to as the UIE-ref dataset, e.g., UIE benchmark datasets UIEB [4] and SUIM-E [12].
(b) Training Procedure: Given the paired UIE-air dataset denoted as $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{N}$, our objective is to train a conditional diffusion model that implies the prior knowledge of mapping transitions from the underwater degradation domain to the real in-air natural domain. We aim to achieve this by minimizing the learning objective $\mathcal{L}_{2}$ (refer to Eq. (9)) over an in-air natural image $x$, conditioned on a synthesized underwater image $y$ (as shown in Fig. 2 (b)). The gradients from the classifier $p(y|x_{t})$ are used directly to guide the reverse process, ensuring the sample $x_{t}$ adhere to the conditioning information $y$. After the training is completed, the prior knowledge of the pre-trained model is utilized to mitigate the adverse impact of unreal synthetic. Subsequently, this model will be guided by additional UIE-ref dataset to cater to specific underwater scenarios, but this fine-tuning or directly continue training may inadvertently compromise the prior knowledge, leading to catastrophic forgetting and mode collapse.
To avoid this issue, thanks to the Classifier Guidance strategy, we can train a classifier to possesses the prior knowledge of the in-air natural domain, distinguishing between in-air natural domain images and underwater domain images. Then, we exploit this classifier along with the additional UIE-ref dataset to jointly fine-tune the diffusion model. This classifier helps retain the prior knowledge throughout the fine-tuning process.
Figure 3: New fine-tuning strategy. From state $x_{T}$, according to Eq. 9, we first adopt a single classifier guidance strategy, using condition $y_{1}$—the input source image—to guide the reverse diffusion process until state $x_{t}$. Then, we switch to the multi-classifier guidance strategy according to Eq. 16. With multi-condition guidance, the intermediate results from $x_{t}$ to $x_{0}$ move towards to the in-air natural domain, mitigating the damage of fine-tuning to the prior knowledge of the pre-trained model.
@@
Figure 4: Illustration of the prompt learning for CLIP-Classifier. (A) Prompt Initialization. Given two text prompts describing the in-air natural image $I_{n}$ and underwater image $I_{u}$. We encode each text to get the initial in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$ and the initial underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$. (B) Prompt Training. We use the cross-entropy loss to constrain the learnable prompts, aligning learnable prompts and images in the CLIP latent space by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$. The base model of CLIP is frozen throughout. The further use of learned prompt for CLIP-Classifier can be seen in Fig. 3.
Stage-two: Fine-tune the Pre-trained Model under Classifier Guidance.
Using Bayes formula, the score function of the multi-condition model under Classifier Guidance can be derived as:
$\bigtriangledown\log p(x_{t}|y_{1},\dots,$ $y_{m})=\bigtriangledown\log(\frac{p(x_{t})p(y_{1},\dots,y_{m})}{p(y_{1},\dots,y_{m})})$ (11)
=$\bigtriangledown$= ▽ $\log p(x_{t})+\bigtriangledown\log p(y_{1},\dots,y_{m}|x_{t}).$
$\bigtriangledown\log p(y_{1},y_{2}|x_{t})=\bigtriangledown\log p(y_{1}|x_{t})+\dots+\bigtriangledown\log p(y_{m}|x_{t}).$ (12)
In our UIE task, there are only two conditions, the source image $y_{1}$ and the in-air natural domain $y_{2}$. So we simplify Eq. (12) by setting $m=2$. Then, substituting this into Eq. (11), we have:
$+\bigtriangledown\log p(y_{2}|x_{t}),$
where $p(y_{1}|x_{t})$ is the same classifier as $p(y|x_{t})$ in Eq. (7) that makes the sample $x_{t}$ adhere to the source image $y_{1}$. The $p(y_{2}|x_{t})$ is the trained classifier that makes the sample $x_{t}$ move towards the in-air natural domain $y_{2}$.
$+(1-\lambda)\bigtriangledown\log p(y_{2}|x_{t}),$
when $\lambda<0.5$, the conditional diffusion model prioritizes the conditioning information $y_{2}$ and moves more towards the in-air natural domain (see in Fig. 3). The value of $\lambda$ is empirically set to 0.4 in our experiments. Meanwhile, according to Eq. (6) and Eq. (9), we rewrite $\epsilon_{\theta}(x_{t},t)$ as $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$, combining the adversarial gradient of the two classifiers:
and the overall learning objective $\mathcal{L}_{3}$ of the multi-guidance diffusion model is depicted in Eq. (16):
$\mathcal{L}_{3}=\min_{\theta}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),y_{1},y_{2},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},y_{1},y_{2},t)\right\|^{2}.$ (16)
This learning objective $\mathcal{L}_{3}$ is directly used in fine-tuning the pre-trained diffusion model. With multi-condition guidance, the intermediate results from $x_{t}$ to $x_{0}$ are constrained by both the source image $y_{1}$ and in-air natural domain $y_{2}$, mitigating the damage that fine-tuning can cause to the prior knowledge of the pre-trained model.
III-B2 Prompt Learning for CLIP-Classifier
@@
In Section III-B1, what we want is to construct a classifier $p(y_{2}|x_{t})$ with conditional information $y_{2}$ (in-air natural domain) and then combine it with the classifier $p(y_{1}|x_{t})$ with conditional information $y_{1}$ (UIE-ref dataset). This combination will jointly control the diffusion model’s generation process, making the result move toward the in-air natural domain. Thanks to the CLIP’s near-perfect text-driven property [43], we design a CLIP-Classifier to implement this vision, and all we need to do is to find prompts that can distinguish between in-air natural images and underwater images. However, prompts for the quality and abstract perception of in-air natural and underwater images often require a lot of manual prompt tuning, and the similar prompts easily lead to huge differences in CLIP score, as shown in Fig. 1. To overcome this issue, CoOp [28] models a prompt’s context words with learnable vectors, extending the application of CLIP-like vision-language models to downstream tasks. For our task, we only need two prompts to characterize in-air natural and underwater images. Therefore, similar to CLIP-LIT [27], we utilize the prompt learning to model a prompt’s text with the longest learnable tensors, maximizing the ability to characterize images. Note that the base model of CLIP is frozen in the entire process. The process of prompt learning involves two key components: prompt initialization and prompt training, as detailed below.
@@
Prompt Initialization. Randomly given two texts describing in-air natural and underwater images respectively, CLIP converts each text into the multi-model embedding space, obtaining the in-air natural image prompt $T_{n}\in\mathbb{R}^{N\times 512}$, and the underwater image prompt $T_{u}\in\mathbb{R}^{N\times 512}$ (refer to Fig. 4 (a)). $N$ represents the length of prompt tokens, with an upper limit of 77. CoOp [28] has experimentally shown that more prompt tokens led to better performance. Thus, we model $T_{n}$ and $T_{u}$ as learnable tensors with $N=77$, maximizing the length of prompt tokens.
@@
Prompt Training. Given an in-air natural image $I_{n}$ and an underwater image $I_{u}$, the image encoder $\Phi(\cdot)$ of the frozen CLIP takes them as the input and extracts image features. On the other hand, by forwarding the prompt $T_{n}$ and the underwater prompt $T_{u}$ to the text encoder $\Psi(\cdot)$, we can obtain the text features. Given a batch of image-text pairs, the prediction probability is calculated as follows:
@@
where $I$ is an image, $I\in\left\{I_{n},I_{u}\right\}$. $cos(\cdot,\cdot)$ denotes cosine similarity. Specially, the $T_{n}/T_{u}$ is shared among all in-air natural/underwater images. Similar to CLIP [43], the prompt learning is trained to align learnable prompts and images in embedding spaces by maximizing the cosine similarity for matched pairs $\left\{T_{n},I_{n}\right\}$ and $\left\{T_{u},I_{u}\right\}$, while minimizing the cosine similarity for incorrect pairs. The cross-entropy loss can be used as a learning objective, and its gradients can be propagated all the way back to update the prompt tensors (see in Fig. 4 (b)). Since there are only two prompts in our prompt engineering, we can simplify the process by directly using the binary cross-entropy to classify in-air natural and underwater images, thereby learning the prompt tensors:
@@
Substituting CLIP-Classifier $\mathcal{L}_{clip}$ into Eq. (15), the new $\epsilon_{\theta}(x_{t},y_{1},y_{2},t)$ is:
@@
where due to the properties of the cosine function, $y_{2}=T_{u}$. Next, we will explore the details of the CLIP-Classifier $\mathcal{L}_{clip}$ guiding the generation process.
@@
III-C New Fine-Tuning Strategy
The forward process of the diffusion model is actually a process in which image semantic information is continuously obscured by noise, always starting from high-frequency information to low-frequency information. In contrast, the reverse generation process smoothly converts an initial Gaussian noise to a realistic image sample through iterative denoising, working from low-frequency information to high-frequency information. For special image-to-image tasks that require consistent content, such as super-resolution [25], image restoration [47], repaint [48], etc., the input image and the corresponding image share low-frequency semantic information in each transition from $x_{t}$ to $x_{t-1}$ during the generation process [49]. Therefore, with the semantic information shared by the input and output images in the low-frequency region, such as the overall structure, image editing and enhancement based on image-to-image diffusion model can start from the high-frequency region instead of the pure Gaussian noise. For instance, SDEdit [50] is to “hijack” the generation process, adding a suitable amount of noise to smooth out high-frequency semantic details, while still preserving the overall structure of the input image, and then this noisy input is used to initialize the SDE to obtain a denoised result that is faithful to the user guidance input.
Obviously, in our case, the underwater degradation images and the reference images share low-frequency semantic information, so we can introduce new conditions at intermediate moments $x_{t}$ to guide the diffusion process (see in Fig. 3). In addition, we further analyze details of how the CLIP-Classifier $\mathcal{L}_{clip}$ guides the generation process. Given a batch of data pairs, we first use Eq. (2) to produce a sequence of intermediate variables $(x_{0},x_{1},\dots,x_{T})$ for each image by perturbing images with Gaussian noise at time $t\in(0,1)$. Then we employ CLIP-Classifier to calculate the classification probability (CLIP score) of the intermediate variables as well as the score difference. As illustrated in Fig. 5 (a), the randomness of the perturbation noise directly causes the CLIP score curves to vibrate violently in small intervals, but the overall trend of the curves is still pronounced. From Fig. 5 (b), we observe that the CLIP-Classifier $\mathcal{L}_{clip}$ mainly works in the discontinuous high-frequency region where $t\in(0,0.10)\cup(0.25,0.50)$, which is consistent with the properties of the image-to-image diffusion model. Based on this finding, we only need to guide the pre-trained diffusion model in the high-frequency region during the fine-tuning process. This new strategy improves the fine-tuning speed compared to the traditional strategy and can increase the speed by up to 10 times, as detailed in Section IV-F.